# Goedel-Prover: A Frontier Model for Open-Source Automated Theorem Proving

### v2

@@ Abstract
We introduce Goedel-Prover, an open-source large language model (LLM) that achieves the state-of-the-art (SOTA) performance in automated formal proof generation for mathematical problems. The key challenge in this field is the scarcity of formalized math statements and proofs, which we tackle in the following ways. We train statement formalizers to translate the natural language math problems from Numina into formal language (Lean 4), creating a dataset of 1.64 million formal statements. LLMs are used to check that the formal statements accurately preserve the content of the original natural language problems. We then iteratively build a large dataset of formal proofs by training a series of provers. Each prover succeeds in proving many statements that the previous ones could not, and these new proofs are added to the training set for the next prover. Despite using only supervised fine-tuning, our final prover significantly outperforms the previous best open-source model, DeepSeek-Prover-V1.5, which employs reinforcement learning. On the miniF2F benchmark, our model achieves a success rate of 57.6% (Pass@32), surpassing DeepSeek-Prover-V1.5 by 7.6%. On PutnamBench, Goedel-Prover successfully solves 7 problems (Pass@512), ranking first on the leaderboard. Furthermore, it generates $29.7$K formal proofs for Lean Workbook problems, nearly doubling the $15.7$K produced by earlier works.
@@ 1 Introduction
Figure 1: (Left) The performance of Pass@32 for whole-proof generation on miniF2F comparing with previous SOTA models. (Middle) A comparison of Goedel-Prover-SFT and DeepSeek-Prover-V1.5 in terms of miniF2F performance across different inference budgets, ranging from Pass@32, 64, 128, …, to $4\times 6400$. (Right) Goedel-Prover-SFT solve 29.7K problems in Lean Workbook. Previously, InternLM2.5-Step-Prover (Wu et al., 2024) and InternLM-Math-Plus (Ying et al., 2024b) collectively solved 15.7K samples.
@@
A significant challenge in training LLMs for theorem proving in formal languages is the scarcity of formalized math statements and proofs. Writing proofs for theorems expressed in formal languages is highly demanding and necessitates considerable domain expertise. Therefore, existing publicly available datasets for formal languages are limited in size. For example, the Lean Workbook (including Lean Workbook Plus) dataset (Ying et al., 2024a; Wu et al., 2024) comprises a total of 140K formal statements, where formal statements refer to problem statements in Lean without proofs. However, only 15.7K of these statements come with formal proofs, which were found by InternLM2.5-StepProver and InternLM-Math-Plus (Ying et al., 2024a; Wu et al., 2024; Ying et al., 2024b). Additionally, the Open Bootstrapped Theorems dataset (Wang et al., 2024) includes 107K statements with proofs sourced from Mathlib44 (mathlib4, 2023). However, we observe a notable difference in the distribution of Mathlib4 compared to that of general problem-solving benchmarks, such as the widely used miniF2F (Zheng et al., 2021). For instance, statements in miniF2F are primarily from high school and require complex reasoning abilities to solve, whereas those in Mathlib4 focus on the simple manipulation of advanced mathematical concepts. (see Figure 7 and 8 in the appendix for a comparison). Furthermore, we find that incorporating Mathlib4 data into training does not consistently improve performance on miniF2F (Section 4).
In contrast to the scarcity of data in formal languages, there is a vast amount of math problems and solutions written in informal language. For example, Numina (Li et al., 2024a) includes 860K high-quality question and answer pairs sourced from MATH (Hendrycks et al., 2021), GSM8K (Cobbe et al., 2021), AMC (aop, ), AIME (MAA, 2024), the AoPS Forum (aop, ), Chinese K-12 Exams (Shao et al., 2024), World Olympiads, and synthetic data (Mitra et al., 2024). We start by training LLMs to formalize the problem statements in these datasets into Lean language. To increase the diversity of the formalization styles, we train two formalizers. One is trained on informal and formal (I-F) statement pairs from Lean Workbook, while the other is trained on I-F statement pairs annotated by Claude-sonnet-3.5 (Anthropic, 2024). We use these two formalizers to formalize the statements and then employ LLMs to ensure that the formal statements preserve the content of the informal statements. Our efforts result in 1.64 million formal statements.
@@
With our large-scale formalized statements, we demonstrate that expert iteration can lead to SOTA performance in whole-proof generation. Specifically,
  • Our model outperforms DeepSeek-Prover-V1.5-RL (the previous SOTA model) by 7.6% on miniF2F, achieving a Pass@32 score of 57.6% compared to DeepSeek-Prover-V1.5-RL’s 50.0%, as shown in Figure 1 (left). It consistently surpasses DeepSeek-Prover-V1.5-RL across all sampling budgets, including Pass@32, 64, and up to 25600, as shown in Figure 1 (middle).
  • We have cumulatively solved 29.7K problems in Lean Workbook, significantly increasing the existing 15.7K proofs found by InternLM2.5-StepProver and InternLM-Math-Plus (Wu et al., 2024; Ying et al., 2024b), as shown in Figure 1 (right).
  • Our model solves 7 problems on PutnamBench by Pass@512111We initially solved 8 problems on PutnamBench. However, after discussing with the authors of PutnamBench, we discovered that one of the problems was mis-formalized. Therefore, this problem is not included in our count, and we report a total of 7 problems here., securing the #1 position on the leaderboard (Table 4).
  • We open source our code [^2] https://github.com/Goedel-LM/Goedel-Prover, model [^3] https://huggingface.co/Goedel-LM/Goedel-Prover-SFT, and the new proofs discovered [^4] https://huggingface.co/datasets/Goedel-LM/Lean-workbook-proofs in the Lean Workbook to facilitate future research.
@@ 3.1 Statement Formalization
Both Formalizer A and B are trained using supervised fine-tuning with Qwen2.5-Coder-32B  [^5] https://huggingface.co/Qwen/Qwen2.5-Coder-32B. The training of these two formalizers takes less than 24 hours on 8 H100 GPUs. Table 1 presents two examples in which both Formalizer A and Formalizer B yield reasonable formalizations. However, our final prover exhibits varying performance on these formalized statements, highlighting the influence of formalization style on model effectiveness.
@@
Example 1 Example 2
Informal Statement The function $f(x)=2^{|x|}+3x^{2}+ax+1$ is an even function, then $a$ equals $a=0$. If $x$ and $\log_{10}x$ are real numbers and $\log_{10}x<0$, show that $-1<x<0$.
Formalizer A Output [Uncaptioned image]
Pass rate: 14/16
Pass rate: 0/16
Formalizer B Output [Uncaptioned image]
Pass rate: 0/16
Pass rate: 5/16
Table 1: Comparison of formalizer outputs for two examples. Comparison of Formalizer Outputs for Two Examples. In Example 1, Formalizer A defines the ”even function” directly by stating $f(-x)=f(x)$. In contrast, Formalizer B first introduces a function called ”IsEven” and then defines the even function using ”IsEven”. Notably, our prover successfully solves the statements provided by Formalizer A but fails with those from Formalizer B. Example 2 is similar; however, our prover fails to solve the statement provided by Formalizer A but succeeds with the one from Formalizer B.
@@
We employ two tests to assess the quality of the formalized statements. First, the formalized statement must conform to Lean syntax and can successfully compile, with the potential proof replaced by the placeholder “:= by sorry”. This syntax check is known as the Compiling Correctness (CC) Test in the literature (Ying et al., 2024a). Second, the formalized statement must accurately capture the original informal problem, incorporating all assumptions, conditions, and implicit definitions. We refer to this second test as the Faithfulness and Completeness (FC) Test. For the FC test, we use Qwen2.5-72B-Instruct [^6] https://huggingface.co/Qwen/Qwen2.5-72B-Instruct with the prompts shown in Figure 4. For each formalized statement, we generate four independent judgments, and the FC score is calculated as #{“Appropriate” in four Judgments}/4. For example, if the four judgments produced by Qwen2.5-72B-Instruct include three “Appropriate” and one “Inappropriate”, the overall FC score is calculated as 0.75. We filter out formalized statements with an FC score less than 0.5.
Figure 4: Prompts for Faithfulness and Completeness (FC) Test.
Model Pass Formalizer A Formalizer B
CC Test Pass@1 76.74% 88.48%
CC Test Pass@8 95.93% 98.59%
FC Test Pass@1 48.06% 80.42%
FC Test Pass@8 88.01% 97.22%
CC + FC Test Pass@1 45.72% 76.41%
CC + FC Test Pass@8 82.33% 95.78%
Table 2: Quality assessment of the formalized statement
@@ 3.2 Expert Iteration
After obtaining a large collection of formalized statements in Section 3.1, we employ expert iteration to train the prover (Liu et al., 2024; Wu et al., 2024; Li et al., 2024b), which is illustrated in Figure 3. Specifically, we first utilize DeepSeek-Prover-V1.5-RL [^7] https://huggingface.co/deepseek-ai/DeepSeek-Prover-V1.5-RL to generate 16 proofs for each statement. We then verify these proofs with the Lean compiler. If at least one proof solves the statement, we retain one proof per statement. In cases where multiple proofs are available, we randomly sample one solution. These collected proofs are used for supervised fine-tuning (SFT) based on DeepSeek-Prover-V1.5-Base [^8] https://huggingface.co/deepseek-ai/DeepSeek-Prover-V1.5-Base, resulting in the iter-1 prover. We continue this expert iteration process; each time, we use the iter-$k$ prover to generate answers and cumulatively collect correct solutions to train DeepSeek-Prover-V1.5-Base for the next iteration, the iter-$(k+1)$ prover. Refer to Appendix A for more details on each iteration.
@@ 4 Results
ProofNet performance.
We observe that the performance on ProofNet exhibits a very low correlation with miniF2F, Lean Workbook and NuminaTest. Specifically, Figure 6 presents the correlation of model performance across the four datasets. The model’s performance on NuminaTest, miniF2F, and Lean Workbook exhibits strong positive correlations, whereas its performance on ProofNet shows a notably low correlation with the other three datasets.
We also observe that the data distribution in ProofNet is significantly different from that of the other three datasets. The problems in ProofNet are primarily drawn from undergraduate pure mathematics textbooks, covering topics such as real and complex analysis, linear algebra, abstract algebra, and topology. These topics largely rely on the abstract and general formulations of mathematical definitions in Mathlib4 (mathlib4, 2023). Other datasets, e.g., miniF2F, largely consist of competition and Olympic-style problems, which require complex reasoning, while only depending on a relatively small set of elementary facts about integers, real numbers, counting, and geometry. We show two examples in Table 5 to illustrate the style difference between ProofNet and miniF2F.
@@ 6 GPT-4o Whole-Proof Generation 1 10
Table 4: Number of problems solved on PutnamBench statements (out of 644). The performance numbers for existing works are taken from the leaderboard. Here $\diamond$ inidicates open-source models.
Figure 6: This figure illustrates the correlation of model performance across four datasets. A lower value indicates a weaker correlation between the model’s performance on two datasets. Notably, ProofNet shows a very low correlation with the other three datasets.
Example from ProofNet Example from miniF2F
Formal Statement [Uncaptioned image] [Uncaptioned image]
Including Mathlib4 in the training data.
Given the close connection between Mathlib4 and ProofNet, we investigate the inclusion of Mathlib4 in the training dataset, following the approaches of Xin et al. (2024b) and Wang et al. (2024). Our results in Table 6 indicate that while adding Mathlib4 enhances the model’s performance on ProofNet, it also leads to a decrease in miniF2F performance. Specifically, we analyze the impact of incorporating Mathlib4 into the training set during the sixth iteration, with the proofs generated by the 5th iteration referred to as “Iter-5 proofs”. We observe that including Mathlib4 increases ProofNet performance from 13.3% to 15.6%, but results in a decline in miniF2F performance from 56.6% to 54.1%, as well as a decrease in NuminaTest performance from 59.2% to 58.8%. We continue to include Mathlib4 in the training dataset from the 6th iteration 6 onward, consistent with DeepSeek-Prover-V1.5-RL (Xin et al., 2024b) and TheoremLamma (Wang et al., 2024). Additional details can be found in the Appendix A.
@@
Table 6: Investigation on ProofNet at the 6th iteration. The results indicate that incorporating Mathlib4 into the training data enhances performance on ProofNet but reduces performance on miniF2F and NuminaTest. We continue to include Mathlib4 in the training dataset from iteration 6 onward, in line with Xin et al. (2024a). Details are in Appendix A.
Proofs found in Lean Workbook.
The Lean Workbook, which includes Lean Workbook-plus (Ying et al., 2024a; Wu et al., 2024), formalizes 140K high-quality problems sourced from AOPS and the Compfiles data. Currently, proofs for only 15.7K statements in Lean Workbook have been found and made open-source by InternLM2.5-StepProver (Wu et al., 2024) and InternLM-Math-Plus (Ying et al., 2024b). In contrast, our model has discovered a significantly larger set of proofs within Lean Workbook, cumulatively solving 29.7K problems. Figure 1 (right) shows a comparison of the number of proofs found by our model and prior works. We open-source all the proofs found by our model to benefit the research community.
Ablation on two formalizers.
In Section 3.1, we discuss that the style of statement formalization can impact the prover’s performance. Therefore, we propose generating statements using two formalizers trained on data from different sources. In this section, we compare a prover trained on a single statement style (using only one formalizer) with a prover trained on both styles (utilizing statements from both Formalizer A and B). We conducted experiments by training the prover at the 8th iteration on statements with proofs collected by the iter-7 prover, and the results are presented in Table 7. The findings indicate that using both formalizers leads to improved performance.
Formalization Model miniF2F ProofNet NuminaTest Average
Formalizer A only 56.5% 13.8% 59.6% 43.3%
Formalizer B only 56.2% 15.2% 60.0% 43.8%
Formalizer A and B 57.6% 15.2% 61.2% 44.7%
Table 7: An ablation study on using two formalizers to formalize the statements. The term ’Formalizer A and B’ refers to using statements formalized by both A and B. In contrast, ’Formalizer A only’ or ’Formalizer B only’ indicates that we include statements formalized exclusively by Formalizer A or B, respectively. Notably, in these experiments, we all add the Mathlib4 training data.

### v3

@@ Abstract
We introduce Goedel-Prover, an open-source language model that achieves state-of-the-art (as of April 5 2025) performance in automated formal proof generation for mathematical problems. A key challenge in this field is the scarcity of formalized mathematical statements and proofs, which we address through the following approaches. First, we train LLMs to convert natural language math problems from the Numina dataset to equivalent formal statements in Lean 4. This process creates the dataset Goedel-Pset-v1, which includes 1.64 million formal statements. Next, we develop a large dataset of formal proofs by training a series of provers. Each new prover can prove many statements that previous ones could not, and these new proofs are added to the training set for the next prover. Finally, we obtain the dataset Goedel-Pset-v1-solved, which contains proofs for over 800K statements from Goedel-Pset-v1. Supervised fine-tuning (SFT) of DeepSeek-Prover-V1.5-Base on Goedel-Pset-v1-solved (i.e., no RL) yields a model that achieves a success rate of 57.6% (Pass@32) on miniF2F benchmark, surpassing the previous leader DeepSeek-Prover-V1.5 (trained using SFT + RL on a proprietary dataset) by 7.6%. On PutnamBench, Goedel-Prover-SFT successfully solves 7 problems (Pass@512), ranking first on the leaderboard. Further RL training (including DPO) improves Goedel-Prover-SFT’s success rate to over 60% (Pass@32) on miniF2F.
To aid future research, we provide extensive discussion of our training methodology and design choices. We also fully open-source our codes, models, and datasets. Additionally, we open-source formal proofs for $29.7$K problems in Lean Workbook, nearly doubling the $15.7$K solved by prior provers.
@@ 1 Introduction
Figure 1: (Left) Pass@32 performance on miniF2F for whole-proof generation, compared to previous SOTA models. (Middle) A comparison of Goedel-Proverand DeepSeek-Prover-V1.5 on miniF2F performance across varying inference budgets, ranging from Pass@32, 64, 128, …, to $4\times 6400$. (Right) Goedel-Prover-SFT solves 29.7K problems in the Lean Workbook. In comparison, InternLM2.5-Step-Prover (Wu et al., 2024) and InternLM-Math-Plus (Ying et al., 2024b) collectively solved 15.7K samples.
@@
A significant challenge in training LLMs for theorem proving in formal languages is the scarcity of formalized math statements and proofs. Writing proofs for theorems expressed in formal languages is highly demanding and necessitates considerable domain expertise. Therefore, existing publicly available datasets for formal languages are limited in size. For example, the Lean Workbook (including Lean Workbook Plus) dataset (Ying et al., 2024a; Wu et al., 2024) comprises a total of 140K formal statements, where formal statements refer to problem statements in Lean without proofs. However, only 15.7K of these statements come with formal proofs, which were found by InternLM2.5-StepProver and InternLM-Math-Plus (Ying et al., 2024a; Wu et al., 2024; Ying et al., 2024b). Additionally, the Open Bootstrapped Theorems dataset (Wang et al., 2024) includes 107K statements with proofs sourced from Mathlib4 (mathlib4, 2023). However, Mathlib4 exhibits significant distribution shift from general problem-solving benchmarks, such as the widely used miniF2F (Zheng et al., 2021). See Section 5 for detail.
In contrast to the scarcity of data in formal languages, there is a vast amount of math problems and solutions written in informal language. For example, Numina (Li et al., 2024a) includes 860K high-quality question and answer pairs sourced from MATH (Hendrycks et al., 2021), GSM8K (Cobbe et al., 2021), AMC (aop, ), AIME (MAA, 2024), the AoPS Forum (aop, ), Chinese K-12 Exams (Shao et al., 2024), World Olympiads, and synthetic data (Mitra et al., 2024). We start by training LLMs to formalize the problem statements in this dataset into Lean. To increase the diversity of the formalization styles, we train two formalizers. One is trained on informal and formal (I-F) statement pairs from Lean Workbook, while the other is trained on I-F statement pairs annotated by Claude-sonnet-3.5 (Anthropic, 2024). We use these two formalizers to formalize the statements and then employ LLMs to ensure that the formal statements preserve the content of the informal statements. Our efforts result in 1.64 million formal statements.
@@
We demonstrate that expert iteration, with large-scale formalized statements can lead to SOTA performance in formal proof generation. Specifically,
To understand the factors behind Goedel-Prover’s strong performance, we provide an in-depth discussion of our training recipe, analyzing the effects of scaling up training data, the diversity introduced by autoformalization, correlations among datasets, and alternative data synthesis strategies. Furthermore, although our final model is trained purely through supervised fine-tuning, we also explore direct preference optimization (DPO) and reinforcement learning (RL) techniques built on top of it. Our Goedel-Prover-DPO and Goedel-Prover-RL achieves a success rate over 60% (Pass@32) on miniF2F. However, we also find that DPO and RL-trained models tend to overfit to “shortcuts” and benefit less from increased inference-time compute.
@@ 3.1 Statement Formalization
Both Formalizer A and B are trained using supervised fine-tuning with Qwen2.5-Coder-32B [^10] https://huggingface.co/Qwen/Qwen2.5-Coder-32B. The training of these two formalizers takes less than 24 hours on 8 H100 GPUs. See Appendix A.1 for examples of formalized statements form two formalizers, where we observe that even for the same problem, the style of the formalized statement can impact the prover’s performance.
@@
We employ two tests to assess the quality of the formalized statements. First, the formalized statement must conform to Lean syntax and can successfully compile, with the potential proof replaced by the placeholder “:= by sorry”. This syntax check is known as the Compiling Correctness (CC) Test in the literature (Ying et al., 2024a). Second, the formalized statement must accurately capture the original informal problem, incorporating all assumptions, conditions, and implicit definitions. We refer to this second test as the Faithfulness and Completeness (FC) Test. For the FC test, we use Qwen2.5-72B-Instruct [^11] https://huggingface.co/Qwen/Qwen2.5-72B-Instruct, details are presented in Appendix A.2.
@@ 3.2 Expert Iteration
After obtaining a large collection of formalized statements in Section 3.1, we employ expert iteration to train the prover (Liu et al., 2024; Wu et al., 2024; Li et al., 2024b), which is illustrated in Figure 3. Specifically, we first utilize DeepSeek-Prover-V1.5-RL [^12] https://huggingface.co/deepseek-ai/DeepSeek-Prover-V1.5-RL to generate 16 proofs for each statement. We then verify these proofs with the Lean compiler. If at least one proof solves the statement, we retain one proof per statement. In cases where multiple proofs are available, we randomly sample one solution. These collected proofs are used for supervised fine-tuning (SFT) based on DeepSeek-Prover-V1.5-Base [^13] https://huggingface.co/deepseek-ai/DeepSeek-Prover-V1.5-Base, resulting in the iter-1 prover. We continue this expert iteration process; each time, we use the iter-$k$ prover to generate answers and cumulatively collect correct solutions to train DeepSeek-Prover-V1.5-Base for the next iteration, the iter-$(k+1)$ prover. Refer to Appendix B for more details on each iteration.
@@ 4 Results
Refer to caption Refer to caption Refer to caption Refer to caption
@@ 6 GPT-4o Whole-Proof Generation 1 10
Table 2: Number of problems solved on PutnamBench statements (out of 644). Goedel-Prover-SFT achieves the first place in the leaderboard. The performance numbers for existing works are taken from the leaderboard. Here $\diamond$ inidicates open-source models.
Proofs found in Lean Workbook.
The Lean Workbook, which includes Lean Workbook-plus (Ying et al., 2024a; Wu et al., 2024), formalizes 140K high-quality problems sourced from AOPS and the Compfiles data. Currently, proofs for only 15.7K statements in Lean Workbook have been found and made open-source by InternLM2.5-StepProver (Wu et al., 2024) and InternLM-Math-Plus (Ying et al., 2024b). In contrast, our model has discovered a significantly larger set of proofs within Lean Workbook, cumulatively solving 29.7K problems, as shown in Figure 1 (right). We open-source all the proofs found by our model to benefit the research community.
5 Dissecting the training recipe
Scaling up the number of formal statements improves model performance.
Figure 5: Model performance under different size of training statement set, illustrating the value of scale.
Figure 5 shows the performance of provers (average on miniF2F, ProofNet and NuminaTest) trained on different sizes of the formal statement set. For each statement, the corresponding proof is obtained using Goedel-Prover-SFT. We observe a consistent improvement in model performance as the size of the statement set increases, underscoring the value of scale in training effective provers.
Increasing the diversity of formalization styles is beneficial.
Table 3 presents the performance of iter-8 provers trained on different formalization styles of statements, with proofs generated by the iter-7 prover. We find that a prover trained on a mixture of styles—combining statements produced by both Formalizer A and Formalizer B—outperforms provers trained on a single formalization style. This result suggests that exposure to diverse formalization patterns improves the model’s generalization and reasoning ability.
Formalization Model miniF2F ProofNet NuminaTest Average
Formalizer A only 56.5% 13.8% 59.6% 43.3%
Formalizer B only 56.2% 15.2% 60.0% 43.8%
Formalizer A and B 57.6% 15.2% 61.2% 44.7%
Table 3: An ablation study on using two formalizers to formalize the statements. Using statements formalized by both formalizers improves the model’s performance, illustrating the value of diverse formalization styles.
Correlations among datasets.
Figure 6: Correlation of model performance across datasets. ProofNet shows low correlation with the others.
We evaluate model performance across different training iterations and hyperparameter settings, and compute the correlation of performance across multiple datasets (see Figure 6). We observe that the model performance on ProofNet is negatively correlated with the performance on miniF2F, Lean Workbook, and NuminaTest. Further more, we investigate the effect of including Mathlib4 in the training data. As shown in Table 4, incorporating Mathlib4 improves performance on ProofNet but leads to a performance drop on miniF2F. These findings suggest a distribution shift between ProofNet/Mathlib4 and the other datasets. Specifically, Mathlib4 and ProofNet tend to focus on the manipulation of mathematical concepts, whereas datasets like miniF2F, Lean Workbook, and NuminaTest contain more Olympiad-style problems that emphasize complex reasoning over formal mathematical content. Illustrative examples are provided in Appendix C. Despite the observed distribution shift, we continue to include Mathlib4 in the training set from the sixth iteration onward, following the approach of DeepSeek-Prover-V1.5-RL (Xin et al., 2024b) and TheoremLamma (Wang et al., 2024), with the aim of enhancing the model’s general capability across a broader range of mathematical domains. Additional training details can be found in Appendix B.
@@ 5 Dissecting the training recipe
Table 4: Incorporating Mathlib4 into the training data enhances performance on ProofNet but reduces performance on miniF2F and NuminaTest, suggesting distribution shift between Mathlib4/ProofNet and other datasets.
Alternative approach for data synthesis.
In addition to autoformalizing statements and use the prover to provide proofs, we also explored alternative strategies for constructing training datasets, focusing on solving difficult problems by a divide-and-conquer strategy. Inspired by Jiang et al. (2022), we implemented the following pipeline: (1) generate a proof for a formal statement using OpenAI’s o1-preview model, (2) extract a high-level “sketch” of the proof and (3) apply DeepSeek-Prover-V1.5-RL to prove the subgoals provided by the sketch. If all the subgoals are successfully completed, we obtain a valid proof for the original problem. Implementation details are provided in Appendix D. However, this pipeline turned out to be ineffective in practice. When applied to the miniF2F validation set (244 problems), it successfully solved only 76 problems—considerably fewer than the 158 problems solvable by DeepSeek-Prover-V1.5-RL alone. Moreover, out of the 76 problems solved, only one is not solved by DeepSeek-Prover-V1.5-RL, implying that the marginal gain from this pipeline is limited.
Exploring DPO and RL training.
We further explored DPO and RL training on top of Goedel-Prover-SFT. We implemented offline Direct Preference Optimization (DPO) (Rafailov et al., 2023) and online Group Relative Policy Optimization (GRPO) (Shao et al., 2024), implementation details are provided in Appendix E. Table 5 shows that although DPO and GRPO improve the model’s Pass@32 performance, the average proof length grows substantially, and the frequency of certain patterns increases sharply. This phenomenon indicates that the model is overfitting to some syntactic patterns or “shortcuts”, which is related to “reward hacking” (Chen et al., 2024). For example, the Lean tactic try allows trying a tactic and continue execution regardless of whether it works or not. Although often harmless—and occasionally useful—its overuse can lead to ineffective proofs and substantial verification costs. The RL-trained model begins to excessively favor this pattern, ultimately impairing its reasoning and generalization capabilities.
Training method
Pass@32
Pass@3200
Average
proof length
Average number of
tactic “try
SFT 57.5% 62.7% 298 1.50
DPO 60.3% 64.6% 486 10.89
Length-penalized DPO 59.8% 63.1% 308 1.11
GRPO 60.5% 63.1% 355 5.16
Table 5: Models’ behavior under different training methods. RL methods show improvement on miniF2F at Pass@32, but the improvement at Pass@3200 is limited . Furthermore, RL models are prone to excessively favor patterns such as try, which also causes the proof length to increase.
Further experiments show that adding a length penalty during DPO training helps reduce this overfitting. However, we observe that scaling up inference-time compute yields significantly smaller gains for models fine-tuned with either GRPO or length-penalized DPO, compared to the SFT model. As shown in Table 5, these models achieve a 3% improvement over Goedel-Prover-SFT on Pass@32, but this gain diminishes when increasing inference-time compute—for example, at Pass@3200. This indicates that RL training may reduce output diversity, leading to less efficient inference-time scaling.